OpenStudy (anonymous):

The equation |4x+2|+3=w has no solution

OpenStudy (zzr0ck3r):

\(|4x+2|+3=w \iff |4x+2|=w-3\) this has no solutions if the right hand side is negative, i.e. \(w-3<0\iff w<3\)

OpenStudy (wolf1728):

Jecka if it has no solution then you have no question correct? Anyway that equation has 2 unknowns so you would need a second equation to find out what x and w specifically equal. Even so, we can see that if x = -2 then w equals 9 and it also equals 13 and so on.